### v1

# Multiplicative lattices with absorbing factorization

Abstract.
In [22], Yassine et al. introduced the notion of $1$-absorbing prime ideals in commutative rings with nonzero identity. In this article, we examine the concept of $1$-absorbing prime elements in C-lattices. We investigate the C-lattices in which every element is a finite product of $1$-absorbing prime elements (we denote them as OAFLs for short). Moreover, we study C-lattices having $2$-absorbing factorization (we denote them as TAFLs for short).
Key words and phrases:
$1$-absorbing prime element, $2$-absorbing element, C-lattice, principally generated lattice.
2010 Mathematics Subject Classification:
06F10, 06F05, 13A15
1. Introduction
Let $L$ be a set together with an inner binary operation $\cdot$ on $L$ and a partial order $\leq$ on $L$ such that $(L,\cdot)$ is a monoid (i.e., $(L,\cdot)$ a commutative semigroup with identity) and $(L,\leq)$ is a complete lattice (i.e., each subset of $L$ has both a supremum and an infimum with respect to $\leq$). For each subset $E\subseteq L$, we let $\bigvee E$ denote the supremum of $E$, called the join of $E$ and we let $\bigwedge E$ denote the infimum of $E$, called the meet of $E$. For elements $a,b\in L$, let $a\vee b=\bigvee\{a,b\}$ and $a\wedge b=\bigwedge\{a,b\}$. Moreover, set $1=\bigvee L$ and set $0=\bigwedge L$. We say that $(L,\cdot,\leq)$ is a multiplicative lattice if for all $x\in L$ and $E\subseteq L$, it follows that $1x=x$ and $x\bigvee E=\bigvee\{xe\mid e\in E\}$.
We recall a few important situations in which multiplicative lattices occur. In what follows, we use the definitions of star operations, ideal systems and the specific star operations/ideal systems $v$, $t$ and $w$ without further mention. For more information on star operations see [12] and for more information on ideal systems see [14]. A profound introduction and study of the $w$-operation can be found in [21].
  • It is well-known that if $R$ is a commutative ring with identity, $L$ is the set of ideals of $R$ and $\cdot:L\times L\rightarrow L$ is the ideal multiplication on $L$, then $(L,\cdot,\subseteq)$ is a multiplicative lattice.
  • Let $D$ be an integral domain and let $\ast$ be a star operation on $D$. Let $L$ be the set of $\ast$-ideals of $D$ together with the $\ast$-multiplication $\cdot_{\ast}:L\times L\rightarrow L$. Then $(L,\cdot_{\ast},\subseteq)$ is a multiplicative lattice.
  • Let $H$ be a commutative cancellative monoid and let $r$ be an ideal system on $H$. Let $L$ be the set of $r$-ideals of $H$ and let $\cdot_{r}:L\times L\rightarrow L$ be the $r$-multiplication. Then $(L,\cdot_{r},\subseteq)$ is a multiplicative lattice.
Let $L$ be a multiplicative lattice and let $e\in L$. For $a,b\in L$, we set $(a:b)=\bigvee\{x\in L\mid xb\leq a\}$. Then $e$ is called weak meet principal if $a\wedge e=(a:e)e$ for each $a\in L$ and $e$ is called weak join principal if $(be:e)=(0:e)\vee b$ for each $b\in L$. Furthermore, $e$ is said to be meet principal if $a\wedge be=((a:e)\wedge b)e$ for all $a,b\in L$ and $e$ is said to be join principal if $((a\vee be):e)=(a:e)\vee b$ for all $a,b\in L$. We say that $e$ is weak principal if $e$ is both weak meet principal and weak join principal. Finally, $e$ is said to be principal ([9]) if $e$ is both meet principal and join principal. An element $a\in L$ is said to be compact if for each subset $F\subseteq L$ with $a\leq\bigvee F$, it follows that $a\leq\bigvee E$ for some finite subset $E$ of $F$. A subset $C\subseteq L$ is called multiplicatively closed if $1\in C$ and $xy\in C$ for each $x,y\in C$. A multiplicative lattice $L$ is called a C-lattice if $L$ is generated under joins by a multiplicatively closed subset $C$ of compact elements. Note that a finite product of compact elements in a C-lattice is again compact. By $L_{*}$ we denote the set of all compact elements of $L$. We say that $L$ is principally generated if every element of $L$ is the join of a set of principal elements of $L$. It is well-known (see [3, Theorem 1.3]) that each principal element of a C-lattice is compact. Moreover, $L$ is said to be join-principally generated if each element of $L$ is the join of a set of join principal elements of $L$. Additionally, a lattice $L$ is called a principal element lattice if every element in $L$ is principal [4].
Let $R$ be a commutative ring with identity, let $D$ be an integral domain, let $H$ be a commutative cancellative monoid, let $\ast$ be a star operation on $D$ and let $r$ be an ideal system on $H$. Note that the lattice of ideals of $R$ is a principally generated C-lattice. The lattice of $\ast$-ideals of $D$ is a C-lattice if and only if $\ast$ is a star operation of finite type. In analogy, it follows that the lattice of $r$-ideals of $H$ is a C-lattice if and only if $r$ is a finitary ideal system. Observe that the lattice of $v$-ideals of $D$ (or of $H$) can fail to be a C-lattice. Also note that even if $\ast$ is of finite type (resp. $r$ is finitary), then the lattice of $\ast$-ideals of $D$ (resp. the lattice of $r$-ideals of $H$) need not be (join-)principally generated. For instance, the $t$-operation is of finite type (resp. the $t$-system is finitary), but the lattice of $t$-ideals is (in general) not (join-)principally generated. We also want to emphasize that the lattice of $w$-ideals of $D$ (resp. of $H$) is a principally generated C-lattice.
An element $a\in L$ is said to be proper if $a<1$, it is called nilpotent if $a^{n}=0$ for some $n\in\mathbb{N}$ and it is called comparable if $a\leq b$ or $b\leq a$ for each $b\in L$. For each $a\in L$, $L/a=\{b\in L\mid a\leq b\}$ is a multiplicative lattice with the multiplication $c\circ d=(cd)\vee a$ for elements $c,d\in L/a$. A proper element $p\in L$ is called prime if $ab\leq p$ implies $a\leq p$ or $b\leq p$ for all $a,b\in L$. A proper element $m\in L$ is said to be maximal in $L$ if for each $x\in L$, $m<x\leq 1$ implies $x=1$. One can easily see that maximal elements are prime. For each $a\in L$, let $\min(a)$ be the set of prime elements of $L$ that are minimal above $a$. The lattice $L$ is called a lattice domain if 00 is a prime element. ${\rm J}(L)$ is defined as the meet of all maximal elements of $L$. For $a\in L$, we define $\sqrt{a}=\bigwedge\{p\in L\mid p$ is prime and $a\leq p\}$. Note that in a C-lattice $L$, $\sqrt{a}=\bigwedge\{p\in L\mid a\leq p$ is a minimal prime over $a\}=\bigvee\{x\in L_{*}\mid x^{n}\leq a$ for some $n\in\mathbb{N}\}$. A proper element $q\in L$ is called primary if $ab\leq q$ implies $a\leq q$ or $b\leq\sqrt{q}$ for every $a,b\in L$. It is well-known that C-lattices can be localized at arbitrary multiplicatively closed subsets $S$ of compact elements as follows. The localization of $a\in L$ at $S$ is defined as $a_{S}=\bigvee\{x\in L\mid xs\leq a$ for some $s\in S\}$. The multiplication on $L_{S}=\{a_{S}\mid a\in L\}$ is defined by $a\circ_{S}b=(ab)_{S}$ for all $a,b\in L_{S}$. Let $p\in L$ be a prime element and $S=\{x\in L_{*}\mid x\nleq p\}$. Then the set $S$ is a multiplicatively closed subset of $L$. In this case, the localization $L_{S}$ is denoted by $L_{p}$. It is well-known that $(L_{p})_{*}=\{a_{p}\in L_{p}\mid a\in L_{*}\}$. Using this, it can be shown that if $L$ is a (principally generated) C-lattice, then $L_{p}$ is also a (principally generated) C-lattice for any prime element $p\in L$ (see [2, Theorem 2.9]). It can also be proved that in a C-lattice $L$, for all $a,b\in L$, $(ab)_{m}=(a_{m}b_{m})_{m}$ for each maximal element $m\in L$ and also, $a=b$ if and only if $a_{n}=b_{n}$ for all maximal elements $n\in L$. For more information on localization, see [2, 3, 9, 17].
In [22], the authors introduced the concept of $1$-absorbing prime ideals in commutative rings with identity. These ideals are generalizations of prime ideals and many authors studied them from different points of view (see [8]). The first aim of this paper is to study $1$-absorbing prime elements in C-lattices. Another (well-known) generalization of $1$-absorbing prime ideals are $2$-absorbing ideals. They have first been mentioned in [7] and in [18], the authors introduced $2$-absorbing elements in multiplicative lattices.
The aforementioned concepts are part of the more general definition, namely that of $n$-absorbing ideals. These types of ideals were introduced and studied by Anderson and Badawi (see [6]). It turns out that $n$-absorbing ideals are not just interesting objects in multiplicative ideal theory, but also in factorization theory. For instance, there is an important connection between $n$-absorbing ideals and the $\omega$-invariant in factorization theory (see [6]). For a profound discussion of the $\omega$-invariant, we refer to [13].
We want to emphasize that the commutative rings in which each ideal is a finite product of $1$-absorbing prime ideals (resp. $2$-absorbing ideals, resp. $n$-absorbing ideals) have already been studied (see [1, 11, 18]). The main goal of this paper is to consider principally generated C-lattices in which various types of elements can be written as finite products of $1$-absorbing prime elements or $2$-absorbing elements.
We continue with a few more basic definitions that will be needed in the sequel. $L$ is said to be a field if $L=\{0,1\}$ and $L$ is called a quasi-local lattice if $1$ is compact and $L$ has a unique maximal element. The dimension of $L$, denoted by $\dim(L)$, is defined to be $\sup\{n\in\mathbb{N}\mid$ there exists a strict chain of prime elements of $L$ of length $n\}$. If $\dim(L)=0$, then $L$ is said to be a zero-dimensional lattice. Note that $L$ is a zero-dimensional lattice if and only if every prime element of $L$ is maximal. We say that a multiplicative lattice is Noetherian if every element of $L$ is compact (see [16, page 352]). A multiplicative lattice is said to be Prüfer lattice if every compact element of $L$ is principal. (For more information about Prüfer lattices, see [3, Theorem 3.4].) A ZPI-lattice is a multiplicative lattice in which every element is a finite product of prime elements [15]. A multiplicative lattice $L$ is said to be a Q-lattice if every element is a finite product of primary elements [20]. A principally generated lattice domain $L$ is called unique factorization lattice domain if every principal element of $L$ is a finite product of principal prime elements.
Our paper is organized as follows. In Section 2, we study the concept of $1$-absorbing prime elements (OA-elements). The relationships among prime elements, primary elements, and TA-elements are studied in Examples 2.2 and 2.3. Propositions 2.7 and 2.8, along with Corollary 2.9, demonstrate that the concepts of prime elements and OA-elements coincide in C-lattices that are not quasi-local. In Section 3, we study C-lattices whose elements have a TA-factorization. We call a C-lattices a TA-factorization lattice (abbreviated as TAFL) if every element possesses a TA-factorization. In Proposition 3.4, we get that $\dim(L)\leq 1$ if $L$ is a principally generated TAFL. In Theorem 3.5, we obtain that a TAFL is ZPI-lattice domain if it is a Prüfer lattice domain. Then, we study the factorization of C-lattices by assuming that all compact elements of $L$ have a factorization into TA-elements, denoted by CTAFL. Finally, we explore the factorization of C-lattices by assuming that all principal elements of $L$ have a factorization into TA-elements, denoted by PTAFL. In Theorem 3.11, we have that if $(L,m)$ is a quasi-local principally generated C-lattice domain, then $L$ is a TAFL if and only if $L$ is a PTAFL and $\dim(L)\leq 1$. In Section 4, we study the factorization of $L$ with respect to the OA-element concept, similar to Section 3. We study C-lattices as a OA-factorization lattice (abbreviated as OAFL) if every element possesses an OA-factorization. Then, we examine the factorization of C-lattices by assuming that all compact elements of $L$ have a factorization into OA-elements, denoted by COAFL. Finally, we explore the factorization of C-lattices by assuming that all principal elements of $L$ have a factorization into OA-elements, denoted by POAFL. Among the many results, in Theorem 4.13, we characterize OAFL, COAFL and lattices which of the join of any two principal elements has an OA-factorization. In Theorem 4.13, we also see that if $L$ is an OAFL, then it satisfies one of the following conditions.
  1. i.
    $L$ is a ZPI-lattice.
  2. ii.
    $L$ is a quasi-local lattice, $m^{2}$ is comparable and $m$ is a nilpotent element.
  3. iii.
    $L$ is a quasi-local lattice domain, $m^{2}$ is comparable and $\bigwedge_{n\in\mathbb{N}}m^{n}=0$.
In Theorem 4.14, we conclude that the following statements are equivalent: $L$ is a ZPI-lattice if and only if $L$ is a Prüfer OAFL if and only if $L$ is a Prüfer POAFL. Theorem 4.15 establishes relationships among the concepts of OAFL, COAFL, TAFL and CTAFL.
2. On 1-absorbing prime elements of C-lattices
Definition 2.1.
Let $L$ be a C-lattice. A proper element $x\in L$ is called a $1$-absorbing prime element or an OA-element if for all $a,b,c\in L\setminus\{1\}$, $abc\leq x$ implies that $ab\leq x$ or $c\leq x$.
It follows immediately from the definition that every OA-element is both a TA-element and a primary element. Moreover, every prime element is an OA-element. We infer that the class of OA-elements of $L$ lies between the classes of prime elements and TA-elements and also between the classes of prime elements and primary elements.
Let $L$ be a C-lattice and let $a\in L$. We obtain the following irreversible right arrows:
  1. (1)
    $a$ is a prime element $\Rightarrow$ $a$ is an OA-element $\Rightarrow$ $a$ is a primary element.
  2. (2)
    $a$ is a prime element $\Rightarrow$ $a$ is an OA-element $\Rightarrow$ $a$ is a TA-element.
We give some examples to show that these arrows are not reversible.
Example 2.2.
{This example is inspired by [10, Example 7]}. Let $L$ be a C-lattice, which having underlying set $\{0,1,a,b,c,d\}$ ordered by $a\leq b\leq d$ and $a\leq c\leq d$, with multiplication $xy=a$ for all $x,y\in\{a,b,c,d\}$. The prime elements of $L$ are 00 and $d$. Moreover, $L$ is a quasi-local lattice. Note that $b$ is an OA-element of $L$ that is not a prime element. In particular, $b$ is a primary TA-element of $L$.
Example 2.3.
We demonstrate that, in general, neither TA-elements nor primary elements are OA-elements. Let ${\rm I}(\mathbb{Z})$ be the lattice of ideals of $\mathbb{Z}$. Note that $(15)$ is a TA-element of ${\rm I}(\mathbb{Z})$ that is not an OA-element of ${\rm I}(\mathbb{Z})$. Furthermore, $(8)$ is a primary element of ${\rm I}(\mathbb{Z})$ that fails to be an OA-element of ${\rm I}(\mathbb{Z})$.
Lemma 2.4.
Let $L$ be a C-lattice. An element $x\in L$ is an OA-element if and only if for all $a,b,c\in L_{*}\setminus\{1\}$, $abc\leq x$ implies that $ab\leq x$ or $c\leq x$.
Proof.
($\Rightarrow$) This is clear.
($\Leftarrow$) Let $abc\leq x$ and $ab\nleq x$ for some $a,b,c\in L\setminus\{1\}$. We show that $c\leq x$ to complete the proof. Since $abc\leq x$, then $a^{\prime}b^{\prime}c^{\prime}\leq x$ for all compact elements $a{\prime},b^{\prime},c^{\prime}\in L$ with $a^{\prime}\leq a$, $b^{\prime}\leq b$ and $c^{\prime}\leq c$. Since $ab\nleq x$, then there are some compact elements $a_{1},b_{1}\in L$ such that $a_{1}\leq a$, $b_{1}\leq b$ and $a_{1}b_{1}\nleq x$. Let $a_{2}=a^{\prime}\vee a_{1}$ and $b_{2}=b^{\prime}\vee b_{1}$. It is clear that $a_{2}$ and $b_{2}$ are compact. Obviously, there is a compact element $c^{*}\in L$ with $c^{*}\leq c$. Note that $(a^{\prime}\vee a_{1})(b^{\prime}\vee b_{1})c^{*}\leq x$ and $(a^{\prime}\vee a_{1})(b^{\prime}\vee b_{1})\nleq x$. We obtain that $c^{*}\leq x$, and thus $c\leq x$. Therefore, $x$ is an OA-element. ∎
Proposition 2.5.
Let $L$ be a C-lattice and let $x\in L$.
  1. (1)
    If $x$ is an OA-element of $L$, then $\sqrt{x}$ is a prime element of $L$ with $(\sqrt{x})^{2}\leq x$.
  2. (2)
    If $x$ is an OA-element of $L$, then $(x:a)$ is a prime element of $L$ for each $a\in L$ with $a\nleq x$.
  3. (3)
    If $(p^{2}:a)\leq x$ for every compact element $a\leq p$, $a\nleq x$ and $x$ is a $p$-primary element of $L$, then $x$ is an OA-element of $L$.
Proof.
(1) Let $x$ be an OA-element of $L$ and let $a,b\in L$ be such that $ab\leq\sqrt{x}$. There is a positive integer $n$ such that $a^{n}b^{n}\leq x$. We can write $a^{m}a^{n-m}b^{n}\leq x$ for a positive integer $m$ with $m<n$. By the assumption, $a^{n}\leq x$ or $b^{n}\leq x$. Then $a\leq\sqrt{x}$ or $b\leq\sqrt{x}$, and thus $\sqrt{x}$ is prime. Now we will show that $(\sqrt{x})^{2}\leq x$. Let $a,b\in L$ be such that $a,b\leq\sqrt{x}$. Then there is an $n\in\mathbb{N}$ with $a^{n}\leq x$. If $n=1$, then we are done. Let $n>2$. Then $a^{n-2}aa\leq x$, and so $a^{2}\leq x$. Similarly, we have that $b^{2}\leq x$. Note that $a(a\vee b)b\leq x$. Then $ab\leq a(a\vee b)\leq x$ or $ab\leq b\leq x$. In any case, we have that $ab\leq x$. Therefore, $(\sqrt{x})^{2}\leq x$.
(2) Let $x$ be an OA-element of $L$ and let $b,c\in L$ be such that $bc\leq(x:a)$. Then $abc\leq x$. By the assumption, $ab\leq x$ or $c\leq x$. Therefore, $b\leq(x:a)$ or $c\leq(x:a)$.
(3) Let $a,b,c\in L_{*}$ be such that $abc\leq x$ and $a\nleq x$. By assumption, $bc\leq\sqrt{x}=p$. Therefore, we obtain that $abc\leq p^{2}$, and thus $bc\leq(p^{2}:a)\leq p$. We infer that $bc\leq x$, and hence $x$ is an OA-element. ∎
Lemma 2.6.
Let $L$ be a C-lattice. If $w\vee u\neq 1$ for some distinct proper elements $u,w\in L$, then $L$ is quasi-local.
Proof.
Let $w\vee u\neq 1$ be distinct proper elements $u,w\in L$. Assume that $L$ is not a quasi-local lattice. There are at least two distinct maximal elements $m_{1},m_{2}\in L$ such that $m_{1}\vee m_{2}=1$, a contradiction. Therefore, $L$ is quasi-local. ∎
Proposition 2.7.
Let $L$ be a C-lattice and let $x\in L$. If $x$ is an OA-element of $L$ that is not prime, then $L$ is quasi-local.
Proof.
Let $x$ be an OA-element of $L$ that is not a prime. By the assumption, $cd\leq x$ for some $c,d\in L$ implies neither $c\leq x$ nor $d\leq x$. If $w\vee u\neq 1$ for each distinct proper elements $w,u\in L$, then we are done by Lemma 2.6. Assume that $w\vee u=1$ for two distinct proper elements $w,u\in L$. Since $wcd\leq x$ and $d\nleq x$, then $wc\leq x$ and similarly, $ucd\leq x$ and $d\nleq x$, then $uc\leq x$. We obtain that $wc\vee uc=(w\vee u)c\leq x$, and hence $c=1c=(w\vee u)c\leq x$, a contradiction. Therefore, $L$ is quasi-local. ∎
Proposition 2.8.
Let $(L,m)$ be a quasi-local C-lattice and let $x\in L$ be proper. Then $x$ is an OA-element if and only if $x$ is a prime element or $m^{2}\leq x<m$.
Proof.
($\Rightarrow$) Without restriction, we can assume that $x$ is not a prime element of $L$. Clearly, there are two proper elements $a,b\in L$ such that $ab\leq x$, $a\nleq x$ and $b\nleq x$. Set $y=m^{2}$. Note that $yab\leq ab\leq x$. Since $a$, $b$ and $y$ are proper elements of $L$ and $b\nleq x$, we have that $ya\leq x$, and hence $mma\leq x$. Moreover, since $a$ and $m$ are proper elements of $L$ and $a\nleq x$, this implies that $m^{2}=mm\leq x$. Since $x$ is not a prime element of $L$, it is obvious that $x<m$.
($\Leftarrow$) If $x$ is a prime element of $L$, then clearly $x$ is an OA-element of $L$. Now let $m^{2}\leq x<m$. Let $a,b,c\in L$ be proper such that $abc\leq x$ and $c\nleq x$. Note that $a\leq m$ and $b\leq m$. We obtain that $ab\leq m^{2}\leq x$. Therefore, $x$ is an OA-element. ∎
As another consequence of Propositions 2.7 and 2.8, we give the following corollary without proof.
Corollary 2.9.
Let $L$ be a C-lattice. Then there is an OA-element of $L$ that is not prime if and only if $L$ is quasi-local with maximal element $m$ such that $m^{2}\neq m$.
Proposition 2.10.
Let $L$ be a principally generated C-lattice and set $m={\rm J}(L)$. The following statements are equivalent.
  1. (1)
    Every proper element of $L$ is an OA-element.
  2. (2)
    Every proper principal element of $L$ is an OA-element.
  3. (3)
    $L$ is quasi-local and $m^{2}=0$.
Proof.
(1) $\Rightarrow$ (2) This is obvious.
(2) $\Rightarrow$ (3) Assume that $L$ is not a quasi-local lattice. Then each proper principal element is a prime element. Note that $L$ is a lattice domain. Let $x\in L$ be a principal element. It follows that $x^{2}$ is a principal prime element. We conclude that $x=x^{2}$, and thus $1=x\vee(0:x)$. Since $x$ is proper, then we have that $x=0$. Consequently, $L$ is field. But this contradicts the fact that $L$ is not a quasi-local lattice. This implies that $L$ is quasi-local with maximal element $m$. We infer that 00 is prime or $m^{2}=0$ by Proposition 2.8. Suppose that $m^{2}\neq 0$. Then there is a nonzero principal element $c\in L$ with $c\leq m^{2}$. By Proposition 2.8, we get that $c^{2}$ is prime element or $m^{2}\leq c^{2}$. If $c^{2}$ is prime, then we have that $c^{2}=c$. Let $m^{2}\leq c^{2}$. We have that $m^{2}\leq c^{2}\leq c\leq m^{2}$, and hence $c^{2}=c$. In any cases, we obtain that $c^{2}=c$, and hence $1=c\vee(0:c)$, since $c$ is principal. Since $L$ is quasi-local, it follows that $c=1$, a contradiction. Therefore, $m^{2}=0$.
(3) $\Rightarrow$ (1) This follows from Proposition 2.8. ∎
Lemma 2.11.
Let $L$ be a join-principally generated C-lattice. If every nonzero element of $L$ is an OA-element, then $\dim(L)=0$.
Proof.
Let $p\in L$ be a prime element and let $m\in L$ a maximal element with $p<m$. Then there is a join principal element $a\in L$ with $a<m$ and $a\nleq p$. Observe that $a$ is a nonzero element that is not nilpotent. By assumption, $a$ is an OA-element of $L$. We have that $a^{3}=a$ or $a^{3}=a^{2}$, since $0\neq a^{3}$ is an OA-element of $L$. If $a^{3}=a$, then it follows that $1=(0:a)\vee a^{2}$.
Since $(0:a)\neq 1$, we have that $a^{2}=1$, which implies that $m=1$, a contradiction. We get a similar result when assuming that $a^{3}=a^{2}$. We conclude that $\dim(L)=0$. ∎
Proposition 2.12.
Let $L$ be a principally generated C-lattice. If $L$ is quasi-local with maximal element $m$ such that $m^{2}$ is comparable, then the following statements are equivalent.
  1. (1)
    Each two principal elements $x,y\in L$ with $m^{2}\leq x$ and $m^{2}\leq y$ imply that $x\leq y$ or $y\leq x$,
  2. (2)
    If $a$ is an OA-element of $L$, then $a$ is prime or $a=m^{2}$.
Proof.
(1) $\Rightarrow$ (2) Let each two principal elements $x,y\in L$ with $m^{2}\leq x$ and $m^{2}\leq y$ satisfy $x\leq y$ or $y\leq x$. Let $a$ be an OA-element of $L$. Suppose that $a$ is not a prime. By Proposition 2.8, we get $m^{2}\leq a<m$. Let $m^{2}<a$. Clearly, there are two principal elements $c,d\in L$ such that $c\leq x$, $c\nleq m^{2}$ and $d<m$, $d\nleq a$. Note that $c,d\nleq m^{2}$. By the assumption, we have that $m^{2}\leq c,d$. Consequently, $c$ and $d$ are OA-elements. Since $d\nleq c$, then $c\leq d$. Therefore, there is an element $v\in L$ with $c=vd$ and hence, we deduce $c\leq m^{2}$, a contradiction. It must be the case that $a=m^{2}$.
(2) $\Rightarrow$ (1) This is clear. ∎
Although we do not derive the result that the meet of two prime elements or two OA-elements yields an OA-element, we deduce the following result.
Lemma 2.13.
Let $L$ be a C-lattice and let $x,y\in L$ be OA-elements that are not prime. Then $x\wedge y$ and $x\vee y$ are OA-elements.
Proof.
Let $x,y\in L$ be OA-elements that are not prime. Then $L$ is quasi-local. By the assumption, we have that $m^{2}\leq x,y$, and hence $m^{2}\leq x\wedge y$ and $m^{2}\leq x\vee y\neq 1$. Since $m^{2}\leq x\wedge y\leq x\vee y\leq m$ and $L$ is quasi-local, then $x\wedge y$ and $x\vee y$ are OA-elements. ∎
Now, we give a relation between OA-elements and lattice domains.
Proposition 2.14.
Let $L$ be a C-lattice. Then 00 is an OA-element of $L$ if and only if $L$ is a lattice domain or $L$ is quasi-local with maximal element $m$ such that $m^{2}=0$.
Proof.
($\Rightarrow$) Let 00 be an OA-element of $L$. Let $L$ be not a lattice domain. Then 00 is not prime, and thus $L$ is quasi-local with maximal element $m$. We infer that $m^{2}=0$.
($\Leftarrow$) This is obvious. ∎
Proposition 2.15.
Let $L$ be a C-lattice. Every TA-element of $L$ is an OA-element of $L$ if and only if the following conditions hold.
  1. (a)
    For each two prime elements $p,q\in L$, we have that $p\leq q$ or $q\leq p$. In particular, $L$ is quasi-local.
  2. (b)
    If $x$ is a TA-element of $L$ and $p\in\min(x)$, then $x=p$ or $p=m$.
Proof.
($\Rightarrow$) Let $p$ and $q$ be prime elements of $L$. Then $p\wedge q$ is a TA-element of $L$. By assumption, $p\wedge q$ is an OA-element. Then by Proposition 2.5(1), we have that $\sqrt{p\wedge q}=p\wedge q$ is prime, and hence $p\wedge q=p$ or $p\wedge q=q$ by [18, Lemma 7]. We obtain that $p\leq q$ or $q\leq p$. Therefore, $L$ is quasi-local. Now, assume that $p$ is minimal prime over $x$. If $x$ is prime, then it is clear that $x=p$. Let $x$ be not a prime element. Then $L$ is quasi-local with maximal element $m$. Then $m^{2}\leq x<p<m$, and thus $p=m$.
($\Leftarrow$) Suppose that $L$ satisfies (a) and (b). Let $x$ be a TA-element of $L$ and let $p$ be a minimal prime element over $x$. By [18, Theorem 3(1)], $p^{2}\leq x\leq p$. If $x=p$, then $x$ is prime, and thus it is clearly an OA-element. Now let $p=m$. Then $m^{2}\leq x\leq m$, and hence $x$ is an OA-element. ∎
Remark 2.16.
Let $L$ be a C-lattice, let $x\in L$ be an OA-element and let $p\in L$ be a prime element of $L$ such that $p\leq x$. Then $x$ is an OA-element of $L/p$.
Proof.
Let $a\circ b\circ c\leq x$ for some $a,b,c\in L/p$. Then $abc\leq x$. By assumption, $ab\leq x$ or $c\leq x$, and hence $a\circ b\leq x$ or $c\leq x$. Consequently, $x$ is an OA-element of $L/p$. ∎
Remark 2.17.
Let $L$ be a C-lattice and let $p\in L$ be a prime element. If $x\in L$ is an OA-element such that $x\leq p$, then $x_{p}$ is an OA-element of $L_{p}$.
Proof.
Let $x\in L$ be an OA-element such that $x\leq p$. Clearly, $x_{p}$ is a proper element of $L_{p}$. Let $a,b,c\in L_{*}$ be such that $a_{p}b_{p}c_{p}\leq x_{p}$. Then $abc\leq x_{p}$, and hence $dabc\leq x$ for some $d\nleq p$. We have that $dab\leq x$ or $c\leq x$. Note that $d_{p}=1$ by [17]. Then $a_{p}b_{p}\leq x_{p}$ or $c_{p}\leq x_{p}$. Therefore, $x_{p}$ is an OA-element of $L_{p}$. ∎
Theorem 2.18.
Let $L$ be a principally generated C-lattice. Every nonzero proper element of $L$ is an OA-element if and only if $L\cong L_{1}\times L_{2}$ where $L_{1},L_{2}$ are fields or $L$ is quasi-local with maximal element $m$ such that $m=\sqrt{0}$ and $m^{2}\leq x$ for every nonzero proper principal element $x\in L$.
Proof.
($\Rightarrow$) Let every nonzero proper element of $L$ be an OA-element. First let $L$ be quasi-local with maximal element $m$. By assumption, every nonzero proper element of $L$ is a TA-element. Now [18, Theorem 8] completes the proof. Now let $L$ be not quasi-local. Then the concepts of prime elements and OA-elements coincide. Let $m_{1}$ and $m_{2}$ be two distinct maximal elements of $L$. Assume that $m_{1}\wedge m_{2}\neq 0$. By the assumption, $m_{1}\wedge m_{2}$ is prime. It can be shown that $m_{1}=m_{2}$, a contradiction. It follows that $m_{1}\wedge m_{2}=0$, and thus $L\cong L/m_{1}\times L/m_{2}$. Note that $L/m_{1},L/m_{2}$ are fields.
($\Leftarrow$) If $L\cong L_{1}\times L_{2}$, where $L_{1}$ and $L_{2}$ are fields, then each nonzero proper element of $L$ is prime, and hence it is an OA-element. Now let $L$ be quasi-local with maximal element $m$ such that $m=\sqrt{0}$ and $m^{2}\leq x$ for every nonzero proper principal element $x\in L$. Let $y$ be a nonzero proper element of $L$. There is some nonzero principal element $c\in L$ with $c\leq y$. We have that $m^{2}\leq c\leq y$, and thus $y$ is an OA-element of $L$. ∎
Proposition 2.19.
Let $L$ be a principally generated quasi-local Noetherian lattice with maximal element $m$. Then every OA-element is prime if and only if $L$ is field.
Proof.
($\Rightarrow$) Since every OA-element is prime, then we obtain that $m^{2}=m$. Therefore, $m=0$ by [3, Theorem 1.4], and thus $L$ is field.
($\Leftarrow$) This is clear. ∎
3. TAFLs and their generalizations
In this section, we study C-lattices whose elements have a TA-factorization. A TA-factorization of an element $x\in L$ means that $x$ is written as a finite product of TA-elements $(x_{k})_{k=1}^{n}$. (Note that the element $1$ is the empty product.) We say that a C-lattice $L$ is a TA-factorization lattice (abbreviated as TAFL) if every element of $L$ has a TA-factorization.
In this section, we investigate the factorization of elements into TA-elements. Firstly, we study C-lattices whose elements possess a TA-factorization, called TAFLs. Next, we study C-lattices whose compact elements have a factorization into TA-elements. We call them CTAFLs. We also explore the C-lattices whose principal elements have a factorization into TA-elements, called PTAFLs. Clearly, every TAFL is both a CTAFL and a PTAFL.
Example 3.1.
As a simple example, it is clear that each prime element is a TA-element, then every ZPI-lattice is a TAFL. By [19, Example 2.1], we have that the lattice of ideals of $\mathbb{Z}[\sqrt{-7}]$ is not a TAFL.
First, we will present some basic results related to TAFL in the following proposition.
Remark 3.2.
Let $L$ be a TAFL, let $L_{1}$ and $L_{2}$ be C-lattices and let $p\in L$ be a prime element.
  1. (1)
    $\min(x)$ is finite for each $x\in L$.
  2. (2)
    $L_{1}\times L_{2}$ is a TAFL if and only if $L_{1}$ and $L_{2}$ are both TAFLs.
  3. (3)
    $L/p$ is a TAFL.
  4. (4)
    $L_{p}$ is a TAFL.
Proof.
(1) Let $x=\prod_{k=1}^{n}x_{k}$ be a TA-factorization of $x$. By [18, Theorem 3], we have that $\min(x_{i})$ is finite. Then $\min(x)$ is finite, since $\min(x)\subseteq\bigcup_{i=1}^{n}\min(x_{i})$.
(2) It is well-known by [18] that $p_{1}=1_{L_{1}}$ and $p_{2}$ is a TA-element of $L_{2}$ or $p_{2}=1_{L_{2}}$ and $p_{1}$ is a TA-element of $L_{1}$ or $p_{1}$ and $p_{2}$ are prime elements of $L_{1}$ and $L_{2}$, respectively if and only if $(p_{1},p_{2})$ is a TA-element of $L$. The rest now follows easily.
(3) Let $y\in L/p$. By the assumption, $y=\prod_{k=1}^{n}x_{k}$ where $x_{i}$ is a TA-element of $L$ for each $i\in[1,k]$. Note that $x_{i}$ is a TA-element of $L/p$ and $y=(\prod_{k=1}^{n}x_{k})\vee p=\bigcirc_{k=1}^{n}x_{k}$. Consequently, $L/p$ is a TAFL.
(4) Recall that if $x$ is a TA-element of $L$, then $x_{p}$ is a TA-element of $L_{p}$. Let $a\in L_{p}$. Then we have that $a=y_{p}$ for some $y\in L$. By assumption, $y$ has a TA-factorization in $L$, meaning $y$ is the finite product of some TA-elements $(y_{k})_{k=1}^{n}$. We have that $a=y_{p}=(\prod_{k=1}^{n}y_{k})_{p}=\bigcirc_{k=1}^{n}(y_{k})_{p}$. This completes the proof. ∎
Lemma 3.3.
Let $L$ be a C-lattice, let $x\in L$ be proper with $\sqrt{x}\in\max(L)$ and let one of the following conditions be satisfied:
  1. (a)
    $L$ is a TAFL.
  2. (b)
    $L$ is a CTAFL and $x$ is compact.
  3. (c)
    $L$ is a PTAFL and $x$ is principal.
Then $x\leq(\sqrt{x})^{2}$ or $(\sqrt{x})^{2}\leq x$.
Proof.
Let $L$ be a TAFL, let $(\sqrt{x})^{2}\nleq x$ and $\sqrt{x}=m$. By [18, Theorem 3], $x$ is not a TA-element. By assumption, $x=\prod_{i=1}^{n}x_{i}$ where $x_{i}$ is TA-element of $L$ and $n\geq 2$. Since $x\leq x_{i}$ for each $i\in[1,n]$ and $\sqrt{x}=m\in\max(L)$, we have that $\sqrt{x_{i}}=m$ for each $i\in[1,n]$. Consequently, $x\leq m^{2}$.
If $L$ is a CTAFL (resp. a PTAFL) and $x$ is a compact (resp. principal), then this can be shown along the same lines as before. ∎
Proposition 3.4.
Let $L$ be a principally generated TAFL. Then $\dim(L)\leq 1$.
Proof.
Observe that since $\dim(L)=\sup\{\dim(L_{q})\mid q\in L$ is a prime element}}\}}, we can assume without restriction that $L$ is quasi-local with maximal element $m\neq 0$. It remains to show that each nonmaximal prime element of $L$ is a minimal prime element. Let $p\in L$ be a nonmaximal prime element. Since $L$ is principally generated, there is some principal element $y\in L$ such that $y\leq m$ and $y\nleq p$.
Moreover, $p_{q}$ is a prime element of $L_{q}$, $y_{q}$ is a principal element of $L_{q}$ and $q_{q}\in\min((p\vee y)_{q})$. If $p_{q}$ is a minimal prime element of $L_{q}$, then $p$ is a minimal prime element of $L$. For these reasons, we can assume without restriction that $m\in\min(p\vee y)$. Since $L$ is quasi-local, this implies that $\sqrt{p\vee y}=m$. Next we verify the following claims.
Claim 1: $m\neq m^{2}$.
Claim 2: $q\leq m^{2}$ for every prime element $q<m$.
Assume the contrary of claim 1 that $m=m^{2}$. Then $m$ is the only TA-element whose radical is $m$. There is a (join) principal element $y$ with $y<m$ and $y\nleq p$. Clearly, $p\vee y=\prod_{i=1}^{k}x_{i}$ where $k$ is a positive integer and $x_{i}$ is a TA-element of $L$ for each $i\in[1,k]$. Note that $m$ is minimal over $p\vee y$. Then $m^{2}\leq x_{i}$ for each $i\in[1,k]$ by [18, Lemma 5]. Therefore, we get that $m^{2k}=m\leq p\vee y\leq m^{k}=m$ by [18, Theorem 3] and Lemma 3.3. Similarly, we have that $m^{2k}=m\leq p\vee y^{2}\leq m^{k}=m$. This implies that $p\vee y=m=p\vee y^{2}$. Note that $((p\vee y^{2}):y)=(p:y)\vee y=p\vee y$. Then $1=((p\vee y):y)=((p\vee y^{2}):y)=p\vee y$, and hence $1=p\vee y=m$, a contradiction.
To show that the second claim is true, assume that there is a prime element $q<m$ with $q\nleq m^{2}$. Also, there is a principal element $b\in L$ with $b\leq m$ and $b\nleq q$. Note that $b^{n}\nleq q$ for a positive integer $n$. We have that $b\in L$ is a nonzero element that is not nilpotent. Since $L$ is a TAFL and $q\nleq m^{2}$, we have that $q\vee b^{3}$ is a TA-element by Lemma 3.3 and [18, Theorem 3]. It follows that $b^{2}\leq q\vee b^{3}$, since $b^{3}\leq q\vee b^{3}$. Note that $1=((q\vee b^{3}):b^{2})=(q:b^{2})\vee b$. We conclude that $(q:b^{2})=1$ or $b=1$. Therefore, $b^{2}\leq q$ or $b=1$, a contradiction. We infer that every prime element $q\in L$ with $q<m$ satisfies $q\leq m^{2}$.
We will return to the proof of the main part. Since $m\neq m^{2}$, there is a nonzero principal element $c\in L$ with $c\nleq m^{2}$. By claim 2, it follows that $\sqrt{c}=m$. By Lemma 3.3, we get $m^{2}\leq c$. Let $0\neq s\leq p$. We have that $s=\prod_{i=1}^{n}y_{i}$ where $y_{i}$ is a TA-element of $L$. Since $p$ is prime, then $y_{j}\leq p$ for some $j\in[1,n]$. Since $y_{j}\leq p<m^{2}\leq c$, then $y_{j}=c\ell_{j}$ for some $\ell_{j}\in L$ because $c$ is weak meet principal. Note that $\ell_{j}\leq p$ because $c\nleq p$. Consequently, $\ell_{j}\leq p\leq m^{2}\leq c$, and thus $\ell_{j}=ct_{j}$ for some $t_{j}\in L$. Therefore, $y_{j}=c^{2}t_{j}$, and hence $\ell_{j}=ct_{j}\leq y_{j}$, since $y_{j}$ is a TA-element. We obtain that $y_{j}=\ell_{j}$, and so $y_{j}=cy_{j}$. Therefore, $s=sc$. Note that $sm\leq s$. Since $s=sc$, we have that $s=sc\leq sm$, and hence $s=sm$. We conclude that $s=0$ by [3, Theorem 1.4], a contradiction. This implies that $p=0$.  ∎
Theorem 3.5.
If $L$ is a TAFL and a Prüfer lattice domain, then $L$ is a ZPI-lattice domain.
Proof.
Let $L$ be a Prüfer lattice domain such that $L$ is also a TAFL. First we have that $L_{m}$ is a TAFL for each maximal element $m\in L$. Let $m\in L$ be maximal. Since $L$ is a Prüfer lattice domain, then $L_{m}$ is a linearly ordered TAFL. From [18, Theorem 10] recall that if $L$ is a Prüfer lattice domain, then the following statements are equivalent: (1) $p$ is a TA-element, (2) $p$ is a prime element of $L$ or $p=p_{1}^{2}$ is a $p_{1}$-primary element of $L$ or $p=p_{1}\land p_{2}$ where $p_{1}$ and $p_{2}$ are some nonzero prime elements of $L$. By using this characterization, we conclude that every TA-element of $L_{m}$ is a finite product of some prime elements. Therefore, $L_{m}$ is a ZPI-lattice (since it is a TAFL), and hence every element of $L_{m}$ is principal and compact. Moreover, $\dim(L_{m})\leq 1$. Therefore, $\dim(L)\leq 1$. Since $L$ is a TAFL domain and $\dim(L)\leq 1$, every nonzero element of $L$ is contained in only finitely many maximal elements. Now one can show that every element of $L$ is compact (since every nonzero element is locally compact and contained in only finitely many maximal elements). Therefore, every element of $L$ is compact. In particular, $L$ is a principal element lattice domain (since $L$ is a Prüfer lattice domain), and hence it is a ZPI-lattice domain. ∎
Next we study CTAFLs.
Remark 3.6.
Let $L$ be a CTAFL, let $L_{1}$ and $L_{2}$ be C-lattices and let $p\in L$ be a prime element.
  1. (1)
    $\min(x)$ is finite for each $x\in L_{*}$.
  2. (2)
    $L_{1}\times L_{2}$ is a CTAFL if and only if $L_{1}$ and $L_{2}$ are both CTAFLs.
  3. (3)
    If $p$ is compact, then $L/p$ is a CTAFL.
  4. (4)
    $L_{p}$ is a CTAFL.
Proof.
(1) This can be proved along the same lines as in the proof of Remark 3.2.
(2) It follows from [18] that $p_{1}=1_{L_{1}}$ and $p_{2}$ is a TA-element of $L_{2}$ or $p_{2}=1_{L_{2}}$ and $p_{1}$ is a TA-element of $L_{1}$ or $p_{1}$ and $p_{2}$ are prime elements of $L_{1}$ and $L_{2}$, respectively if and only if $(p_{1},p_{2})$ is a TA-element of $L$. The rest is straightforward.
(3) Let $p$ be compact. Observe that every element $a\in L$ with $a\geq p$ is compact in $L$ if and only if $a$ is compact in $L/p$. Now, let $y\in L/p$ be compact. Then $y\geq p$. By the assumption, $y=\prod_{k=1}^{n}x_{k}$ where $x_{i}$ is a TA-element of $L$. Note that $x_{i}$ is a TA-element of $L/p$. Then we get that $y=(\prod_{k=1}^{n}x_{k})\vee p=\bigcirc_{k=1}^{n}x_{k}$. Therefore, $L/p$ is a CTAFL.
(4) This can be shown along similar lines as in Remark 3.2(4). ∎
Proposition 3.7.
Let $L$ be a principally generated CTAFL that satisfies the ascending chain condition on prime elements. Then $\dim(L)\leq 1$.
Proof.
Observe that for each prime element $q\in L$, we have that $L_{q}$ is a principally generated CTAFL that satisfies the ascending chain condition on prime elements. Since $\dim(L)=\sup\{\dim(L_{q})\mid q\in L$ is a prime element}}\}}, we can assume without restriction that $L$ is quasi-local with maximal element $m\neq 0$. It remains to show that each nonmaximal prime element of $L$ is a minimal prime element. Let $p\in L$ be a nonmaximal prime element. Since $L$ is principally generated, there is some principal element $y\in L$ such that $y\leq m$ and $y\nleq p$.
Since $L$ is a C-lattice, there exists some $q\in\min(p\vee y)$ such that $q\leq m$. Clearly, $L_{q}$ is a principally generated CTAFL with maximal element $q_{q}$ that satisfies the ascending chain condition on prime elements. Moreover, $p_{q}$ is a prime element of $L_{q}$, $y_{q}$ is a principal element of $L_{q}$ and $q_{q}\in\min((p\vee y)_{q})$. If $p_{q}$ is a minimal prime element of $L_{q}$, then $p$ is a minimal prime element of $L$. For these reasons, we can assume without restriction that $m\in\min(p\vee y)$. Since $L$ is quasi-local, this implies that $\sqrt{p\vee y}=m$. Next we verify the following claims.
Claim 1: $m\neq m^{2}$.
Claim 2: $q\leq m^{2}$ for every prime element $q\in L$ with $q<m$.
First we prove claim 1. It follows from Remark 3.6(1) that $\min(x)$ is finite for each compact element $x\in L$. Since $L$ satisfies the ascending chain condition on prime elements, it follows from [16, Theorem 2] that $p=\sqrt{d}$ for some compact element $d\in L$. Observe that $m=\sqrt{d\vee y}$ and $d\vee y$ is compact. Consequently, $d\vee y=\prod_{i=1}^{t}x_{i}$ for some positive integer $t$ and some TA-elements $x_{i}\in L$. Clearly, $\sqrt{x_{i}}=m$ for each $i\in[1,t]$, and thus $m^{2}\leq x_{i}\leq m$ by [18, Theorem 3] for each $i\in[1,t]$. Assume to the contrary that $m=m^{2}$. Then $d\vee y=m$, and hence $p\vee y=m$. We infer that $m=m^{2}\leq p\vee y^{2}\leq p\vee y\leq m$, and thus $p\vee y=p\vee y^{2}$. Since $y$ is (join) principal, we obtain that $1=((p\vee y):y)=((p\vee y^{2}):y)=p\vee y=m$, a contradiction. This implies that $m\not=m^{2}$.∎(Claim 1)
Now we prove claim 2. Assume that there is a prime element $q\in L$ such that $q<m$ and $q\nleq m^{2}$. Since $L$ is principally generated, there is a principal element $b\in L$ such that $b\leq m$ and $b\nleq q$. Note that $b^{2}\nleq q$. Since $q\nleq m^{2}$ and $L$ is a C-lattice, there is a compact element $a\in L$ such that $a\leq q$ and $a\nleq m^{2}$. Since $a$ and $b$ are compact, we have that $a\vee b^{3}$ is compact. Note that $a\vee b^{3}$ is a TA-element, since $L$ is a CTAFL and $a\nleq m^{2}$. Since $b^{3}\leq a\vee b^{2}$ and $a\vee b^{2}$ is a TA-element, we get that $b^{2}\leq a\vee b^{3}$. Note that $1=((a\vee b^{3}):b^{2})=(a:b^{2})\vee b$. Since $b\leq m$ and $L$ is quasi-local, we have that $(a:b^{2})=1$. Consequently, $b^{2}\leq a\leq q$, a contradiction.∎(Claim 2)
It is sufficient to show that $p=0$. (Then $p$ is a minimal prime element of $L$ and we are done.) Since $m\neq m^{2}$ by claim 1 and $L$ is principally generated, there is a principal element $c\in L$ such that $c\leq m$ and $c\nleq m^{2}$. By claim 2, we have that $\sqrt{c}=m$. Furthermore, Lemma 3.3 implies that $m^{2}\leq c$. Let $s\in L$ be compact such that $s\leq p$. It follows that $s=\prod_{i=1}^{k}y_{i}$ for some positive integer $k$ and some TA-elements $y_{i}\in L$. Obviously, there is some $j\in[1,k]$ such that $y_{j}\leq p$. Since $y_{j}\leq p\leq m^{2}\leq c$ and $c$ is (weak meet) principal, we infer that $y_{j}=c\ell$ for some $\ell\in L$. Note that $\ell\leq p$, since $c\nleq p$. Consequently, $\ell\leq p\leq m^{2}\leq c$, and thus $\ell=ct$ for some $t\in L$. This implies that $y_{j}=c^{2}t$, and hence $\ell=ct\leq y_{j}$ (since $y_{j}$ is a TA-element of $L$). Therefore, $y_{j}=\ell$, and thus $y_{j}=cy_{j}$. We infer that $s=sc$, and hence $s=sc\leq sm\leq s$. We conclude that $s=sm$. It is an immediate consequence of [3, Theorem 1.4] that $s=0$. Finally, we have that $p=0$ (since $L$ is a C-lattice). ∎
Next we study PTAFLs. We start with a simple observation.
Remark 3.8.
Let $L$ be a Prüfer lattice. Then $L$ is a CTAFL if and only if $L$ is a PTAFL.
Proof.
This is obvious, since every compact element in a Prüfer lattice is principal. ∎
Note that Proposition 3.7 does not hold for PTAFLs. To show that, we consider the following example.
Example 3.9.
Note that if $L$ is the lattice of ideals of a local two-dimensional unique factorization domain $D$ (e.g. take $D=K[X,Y]_{(X,Y)}$ where $K$ is a field and $X$ and $Y$ are indeterminates over $K$), then $L$ is a quasi-local principally generated PTAFL that satisfies the ascending chain condition on prime elements and $\dim(L)=2$.
Theorem 3.10.
Let $(L,m)$ be a quasi-local principally generated C-lattice domain such that $\dim(L)\leq 1$, $m^{2}$ is comparable and $\bigwedge_{n\in\mathbb{N}}m^{n}=0$. Then $L$ is a TAFL.
Proof.
Assume that $m=m^{2}$. Since $L$ is a lattice domain and $\bigwedge_{n\in\mathbb{N}}m^{n}=0$, we infer that $m=0$. Therefore, $L$ is a field and hence we get the desired properties. Now, suppose that $m\neq m^{2}$. Then there is a principal element $x\in L$ with $x\nleq m^{2}$. By the assumption, we get that $m^{2}<x$. Since $\dim(L)\leq 1$, then we conclude that $x$ is a TA-element by [18, Theorem 3]. Since $x$ is meet principal, then we conclude that $m^{2}=xm$. Let $z\in L$ be proper. If $z=0$, then it is clear that it has a TA-factorization under the assumption that $L$ is a lattice domain. Now let $z\not=0$. If $m^{2}\leq z$, then the proof is complete by [18, Theorem 3]. Now let $z\leq m^{2}$. Let $n$ be the largest positive integer satisfying $z\leq m^{n}$. We conclude that $z\leq m^{n}=x^{n-1}m$. Consequently, $z\leq x^{n-1}$. Note that $z=x^{n-1}a$ for some $a\in L$ since $x^{n-1}$ is principal. Suppose that $a\leq m^{2}$. We conclude that $a\leq x$. Moreover, we can write $a=xb$ for some $b\in L$. Consequently, we obtain $z=x^{n}b\leq x^{n}m=m^{n+1}$, leading to a contradiction. This implies that $m^{2}\leq a$, and thus $a$ is a TA-element. Therefore, $z$ has a TA-factorization. ∎
Theorem 3.11.
Let $(L,m)$ be a quasi-local principally generated C-lattice domain. Then $L$ is a TAFL if and only if $\dim(L)\leq 1$ and $L$ is a PTAFL. If these equivalent conditions are satisfied, then $\bigwedge_{n\in\mathbb{N}}m^{n}=0$ and $m^{2}$ is comparable.
Proof.
()($\Rightarrow$)( ⇒ ): This is follows from Proposition 3.4.
()($\Leftarrow$)( ⇐ ): Let $L$ be a PTAFL such that $\dim(L)\leq 1$. First, assume that $m=m^{2}$. Since $\dim(L)\leq 1$, then $m$ is the only TA-element whose radical is $m$. We infer that each proper nonzero element of $L$ is equal to $m$ and we are done.
Now let $m\neq m^{2}$. Then there is a principal element $x\in L$ such that $x\leq m$ and $x\nleq m^{2}$. Since $x\nleq m^{2}$, we infer that $x$ is a TA-element (since $x$ cannot be the product of more than one TA-element). From [18, Theorem 3], we get that $m^{2}\leq x$ since $\sqrt{x}=m$. Since $x$ is a (weak meet) principal element, we conclude that $m^{2}=xm$. Let $z\in L$ be proper. We have to show that $z$ has a TA-factorization. If $z=0$, then we are done, since $L$ is lattice domain. Therefore, we can assume without restriction that $z\neq 0$.
Next we show that $\bigwedge_{n\in\mathbb{N}}m^{n}=0$. Assume the contrary that $\bigwedge_{n\in\mathbb{N}}m^{n}\neq 0$. Note that each nonzero TA-element $v\in L$ satisfies $m^{2}\leq v$ by [18, Lemma 2]. Clearly, there is some nonzero principal element $x\in L$ such that $x\leq\bigwedge_{n\in\mathbb{N}}m^{n}$. We have that $x=\prod_{i=1}^{k}a_{i}$ where $a_{i}$ is a TA-element for each $i\in[1,k]$. We obtain that $(m^{2})^{k}\leq x\leq(m^{4})^{k}\leq(m^{2})^{k}$, and hence $x=x^{2}$. Since $x$ is principal, we infer that $1=x\vee(0:x)$. Therefore, we obtain that $x=1$, a contradiction.
First let $z\leq m^{2}$. Let $n$ be the largest positive integer satisfying $z\leq m^{n}$. We conclude that $z\leq m^{n}=x^{n-1}m$, and thus $z\leq x^{n-1}$. Note that $z=x^{n-1}a$ for some $a\in L$ since $x^{n-1}$ is principal. Suppose that $a\leq m^{2}$. It follows that $a\leq x$. Moreover, we can write $a=xb$ for some $b\in L$. Consequently, we obtain $z=x^{n}b\leq x^{n}m=m^{n+1}$, a contradiction. Let $a\nleq m^{2}$. There is a principal element $a^{\prime}\in L$ such that $a^{\prime}\leq a$ but $a^{\prime}\nleq m^{2}$. Since $a^{\prime}\nleq m^{2}$, we know that $a^{\prime}$ is a TA-element of $L$. Therefore, $m^{2}\leq a^{\prime}\leq a$. We conclude that $a$ is a TA-element. This implies that $z=x^{n-1}a$ has a TA-factorization.
Finally, let $z\nleq m^{2}$. Then there is a principal element $y\in L$ such that $y\leq z$ and $y\nleq m^{2}$. We have that $y$ is a TA-element, and hence $m^{2}\leq y\leq z$. This implies that $z$ is a TA-element. Consequently, $L$ is a TAFL.
It remains to show that $m^{2}$ is comparable. Let $x\in L$ be proper such that $x\nleq m^{2}$. Since $L$ is a TAFL, we conclude that $x$ is a TA-element. From [18, Lemma 2], we get that $m^{2}\leq x$ since $\dim(L)\leq 1$. ∎
Proposition 3.12.
Let $(L,m)$ be a quasi-local principally generated C-lattice such that $L$ is not a lattice domain. Then $L$ is a TAFL if and only if $\dim(L)=0$ and $L$ is a PTAFL.
Proof.
()($\Rightarrow$)( ⇒ ): Let $L$ be a TAFL. By proof of Proposition 3.4, we know that if $p\in L$ is a nonmaximal prime element of $L$, then $p=0$. Therefore, $\dim(L)\leq 1$. Since $L$ is not a lattice domain, we infer that $\dim(L)=0$.
()($\Leftarrow$)( ⇐ ): Let $\dim(L)=0$ and let $L$ be a PTAFL. Note that $m$ is nilpotent element of $L$ since $0=\prod_{i=1}^{k}x_{i}$ where $x_{i}$ is a TA-element with $\sqrt{x_{i}}=m$ for each $i\in[1,k]$, and hence $(m^{2})^{k}\leq 0$. This implies that $0=m^{2k}$, and thus 00 has a TA-factorization. If $m=m^{2}$, we get that $m=0$. Now let $m\neq m^{2}$. Then there is a principal element $z\in L$ such that $z\leq m$ and $z\nleq m^{2}$. Since $z\nleq m^{2}$, we infer that $z$ is a TA-element (since $z$ cannot be the product of more than one TA-element). From [18, Lemma 2], we get that $m^{2}\leq z$ since $\sqrt{z}=m$. Since $z$ is a (weak meet) principal element, we conclude that $m^{2}=zm$. Let $x\in L$ be nonzero. If $m^{2}\leq x$, then $x$ is a TA-element because it is a primary element as shown by [18, Lemma 5].
Next let $x\leq m^{2}$. Let $n$ be the largest positive integer for which $x\leq m^{n}$. We conclude that $x\leq m^{n}=z^{n-1}m$, which implies $x\leq z^{n-1}$. Note that $x=z^{n-1}a$ for some $a\in L$, given that $z^{n-1}$ is a principal element. Assume that $a\leq m^{2}$. This implies that $a\leq z$, and hence $a=zb$ for some $b\in L$. We conclude $x=z^{n}b\leq z^{n}m=m^{n+1}$, a contradiction. Therefore, $a\nleq m^{2}$. There is a principal element $a^{\prime}\in L$ such that $a^{\prime}\leq a$ and $a^{\prime}\nleq m^{2}$. Since $a^{\prime}\nleq m^{2}$, we have that $a^{\prime}$ is a TA-element. Therefore, $m^{2}\leq a^{\prime}\leq a$. Consequently, $a$ is a TA-element. It follows that $x=z^{n-1}a$ has a TA-factorization.
Finally, let $x\nleq m^{2}$. Then there is a principal element $y\in L$ such that $y\nleq m^{2}$ and $y\leq x$. We have that $y$ is a TA-element, and hence $m^{2}\leq y\leq x$. We infer that $x$ is a TA-element. Therefore, $L$ is a TAFL. ∎
Remark 3.13.
Let $(L,m)$ be a quasi-local principal element TAFL domain. Then every proper element of $L$ is a power of $m$.
Proof.
We know that a TA-element of $L$ equals $m$ or $m^{2}$ by [18, Theorem 9]. This completes the proof. ∎
4. OAFLs and their generalizations
In this section, we study the factorization of elements of $L$ with respect to the OA-elements, similar to the previous section. It consists of three parts. The first part involves C-lattices, whose elements possess an OA-factorization, called a OAFLs. Next, we examine C-lattices whose compact elements have a factorization into OA-elements, called a COAFLs. Finally, we explore the C-lattices whose principal elements have a factorization into OA-elements, called POAFLs. It can easily be shown that every OAFL is both a COAFL and a POAFL. We continue by presenting some results related to OAFL.
Example 4.1.
Let $L$ be the lattice of ideals of $\mathbb{Z}[2i]$. Note that $(2+2i)$ has no OA-factorization. Therefore, $L$ is not an OAFL.
Remark 4.2.
Let $L$ be an OAFL and let $p\in L$ be a prime element.
  1. (1)
    $\min(x)$ is finite for each $x\in L$.
  2. (2)
    $L$ is both a Q-lattice and a TAFL.
  3. (3)
    $L_{p}$ is an OAFL.
  4. (4)
    $L/p$ is an OAFL.
Proof.
(1) Let $x\in L$ and let $x=\prod_{k=1}^{n}x_{k}$ be an OA-factorization of $x$ for some OA-elements $x_{i}\in L$. Recall that each OA-element is a TA-element and $\min(a)$ is finite for a TA-element $a\in L$ by [18, Theorem 3]. Hence $\min(x)$ is finite since $\min(x)\subseteq\bigcup_{i=1}^{n}\min(x_{i})$.
(2) Since every OA-element of $L$ is a TA-element and a primary element, we have that $L$ is both a Q-lattice and a TAFL.
(3) Let $y\in L_{p}$ be proper. Then there is a proper element $x\in L$ such that $y=x_{p}$. By the assumption, we get such a factorization $x=\prod_{k=1}^{n}x_{k}$ for some OA-elements $x_{i}\in L$. By Remark 2.17, we know that $(x_{i})_{p}$ is an OA-element of $L_{p}$, and thus $y=x_{p}=(\prod_{k=1}^{n}x_{k})_{p}=\bigcirc_{k=1}^{n}(x_{k})_{p}$. This completes the proof.
(4) Let $y\in L/p$. By the assumption, $y=\prod_{k=1}^{n}x_{k}$ where $x_{i}$ is an OA-element of $L$ for each $i\in[1,k]$. Note that $x_{i}$ is an OA-element of $L/p$. Then we get that $y=(\prod_{k=1}^{n}x_{k})\vee x=\bigcirc_{k=1}^{n}x_{k}$. Observe that $L/p$ is an OAFL. ∎
Corollary 4.3.
Let $L$ be a principally generated OAFL. Then $\dim(L)\leq 1$.
Proof.
This is an immediate consequence of Remark 4.2(2) and Proposition 3.4. ∎
Lemma 4.4.
Let $(L,m)$ be a quasi-local principally generated C-lattice such that $m^{2}$ is comparable and $m$ is nilpotent or $L$ is a lattice domain with $\bigwedge_{n\in\mathbb{N}}m^{n}=0$. Then $L$ is an OAFL and every proper principal element of $L$ is a finite product of principal OA-elements.
Proof.
First let $m=m^{2}$. Since $m$ is nilpotent or $L$ is lattice domain with $\bigwedge_{n\in\mathbb{N}}m^{n}=0$, we have that $m=0$. Therefore, $L$ is a field and hence it satisfies the desired conditions.
Now let that $m\neq m^{2}$. There is a proper principal element $x\in L$ with $x\nleq m^{2}$. By the assumption, we get that $m^{2}<x$. We conclude that $x$ is an OA-element. Since $x$ is meet principal, then we get that $m^{2}=x(m^{2}:x)$. We have that $m^{2}=xm$. Let $z\in L$ be proper. First let $z=0$. If $L$ is lattice domain, then it is clear that $z$ is a principal OA-element. Let $m$ is nilpotent. Then there is a positive integer $n$ with $m^{n}=0$. Note that $z=x^{n}$ is a finite product of principal OA-elements.
Next let $z\not=0$. If $m^{2}<z$, then $z$ is an OA-element and we are done. Now let $z\leq m^{2}$. Let $n$ be the largest positive integer satisfying $z\leq m^{n}$. Therefore, $z\leq m^{n}=x^{n-1}m$, and thus $z\leq x^{n-1}$. Note that $z=x^{n-1}a$ for some $a\in L$ since $x^{n-1}$ is principal. If $z$ is principal, then we can by [5, Theorem 9] assume that $a$ is principal. Suppose that $a\leq m^{2}$. We have that $a\leq x$. Moreover, $a=xb$ for some $b\in L$. Consequently, we obtain $z=x^{n}b\leq x^{n}m=m^{n+1}$, a contradiction. Therefore, $m^{2}\leq a$ and so $a$ is an OA-element by Proposition 2.8. We infer that $z$ has an OA-factorization and if $z$ is principal, then $z$ is a finite product of principal OA-elements. ∎
Next we study COAFLs.
Remark 4.5.
Let $L$ be a COAFL and let $p\in L$ be a prime element.
  1. (1)
    $\min(x)$ is finite for each $x\in L_{*}$.
  2. (2)
    $L$ is a CTAFL.
  3. (3)
    If $p$ is compact, then $L/p$ is a COAFL.
  4. (4)
    Then $L_{p}$ is a COAFL.
Proof.
(1) This can be shown along the lines of the proof of Remark 4.2.
(2) This is clear, since every OA-element of $L$ is a TA-element.
(3) Let $p$ be compact. Note that every element $a\in L$ with $a\geq p$ is compact in $L$ if and only if $a$ is compact in $L/p$. Now, let $y\in L/p$ be compact. Then $y\geq p$. By the assumption, $y=\prod_{k=1}^{n}x_{k}$ where $x_{i}$ is an OA-element of $L$. Note that $x_{i}$ is an OA-element of $L/p$. Then we get that $y=(\prod_{k=1}^{n}x_{k})\vee x=\bigcirc_{k=1}^{n}x_{k}$. Consequently, $L/p$ is a COAFL.
(4) This can proved along similar lines as in Remark 2.17. ∎
Proposition 4.6.
Let $L$ be a quasi-local COAFL. Then each minimal nonmaximal prime element of $L$ is a weak meet principal element.
Proof.
Let $p\in L$ be a minimal nonmaximal prime element of $L$. First we show that $x=p(x:p)$ for each compact element $x\in L$ with $x\leq p$. Let $x\in L$ be a compact element of $L$ with $x\leq p$. By the assumption, we can write $x=\prod_{i=1}^{n}x_{i}$ where $n$ is a positive integer and $x_{i}$ is an OA-element of $L$ for each $i\in[1,n]$. Since $x\leq p$, we have that $x_{i}\leq p$ for some $i\in[1,n]$. If $x_{i}$ is not a prime element, then $m^{2}\leq x_{i}\leq p$ by Proposition 2.8. We infer that $m=p$, a contradiction. Consequently, $x_{i}$ is a prime element, and hence $x_{i}=p$. We have that $x=pz$, where $z=\prod_{k=1,k\not=i}^{n}x_{k}$. Observe that $x=p(x:p)$.
Let $y\in L$ be such that $y\leq p$. Since $L$ is a C-lattice, we have that
$y=\bigvee\{v\in L_{*}\mid v\leq y\}=\bigvee\{p(v:p)\mid v\in L_{*},v\leq y\}=p\bigvee\{(v:p)\mid v\in L_{*},v\leq y\}.$
Set $w=\bigvee\{(v:p)\mid v\in L_{*},v\leq y\}$. Then $w\in L$ and $y=pw$, and thus $p$ is weak meet principal. ∎
Next we study POAFLs. We start with a simple observation.
Remark 4.7.
Let $L$ be a Prüfer lattice. Then $L$ is a COAFL if and only if $L$ is a POAFL.
Proof.
This is obvious, since every compact element in a Prüfer lattice is principal. ∎
Lemma 4.8.
Let $L$ be a quasi-local principally generated POAFL. Then every minimal nonmaximal prime element of $L$ is a principal element.
Proof.
Let $p\in L$ be a minimal nonmaximal prime element of $L$. Let $a\in L$ be a principal element of $L$ with $a\leq p$. By the assumption, we can write $a=\prod_{k=1}^{n}q_{k}$ where $q_{i}$ is an OA-element of $L$ for $i\in[1,n]$. Since $a\leq p$, then $q_{j}\leq p$ for some $j\in[1,n]$. If $q_{j}$ is an OA-element that is not prime, then $m^{2}\leq q_{j}\leq p$. Therefore, $m=p$, a contradiction. Assume that $q_{i}$ is prime. Then we get that $q_{i}=p$. We infer that $a=p\prod_{k=1,k\not=i}^{n}q_{k}$, and thus $a=pb$ where $b=\prod_{k=1,k\not=i}^{n}q_{k}$. Obviously, $a=p(a:p)$. Now, take an element $c\in L$ with $c\leq p$. By assumption, $c=\bigvee A$ for some set $A$ of principal elements of $L$. We conclude that $c=\bigvee A=\bigvee\{p(a:p)\mid a\in A\}=p(\bigvee\{(a:p)\mid a\in A\})$. Therefore, $p$ is a weak meet principal element, and hence $p$ is a principal element by [3, Theorem 1.2]. ∎
Lemma 4.9.
Let $(L,m)$ be a quasi-local principally generated C-lattice. If the join of any two principal elements of $L$ has an OA-factorization, then every nonmaximal prime element of $L$ is a principal element and $\dim(L)\leq 2$.
Proof.
Let the join of any two principal elements of $L$ have an OA-factorization. First, we show that $\dim(L)\leq 2$. Let $p\in L$ be a nonmaximal prime element of $L$. Consider that $L_{p}$ is quasi-local and $L_{p}$ is generated by the set of elements $\{a_{p}\mid a\in L$ is principal}}\}}. We can see that the join of any two principal element of $L_{p}$ has a prime factorization. Because $a_{p}\vee_{p}b_{p}=(a\vee b)_{p}=(\prod_{i=1}^{n}x_{i})_{p}=\bigcirc_{i=1}^{n}(x_{i})_{p}$ where $(x_{i})_{p}$ is a prime element of $L_{p}$. Consequently, $L_{p}$ is a ZPI-lattice by [15, Theorem 8], and thus $\dim(L_{p})\leq 1$. Therefore, $\dim(L)\leq 2$.
Let $q$ be a nonmaximal prime element of $L$. Assume that $qm=q$. If $q\in\min(0)$, then it is clear that $q=0$ by Nakayama’s Lemma since $q$ is principal. Now let $q$ be not minimal. There is a $p\in\min(0)$ such that $p<q$, and thus there is a principal element $c\in L$ such that $c\nleq p$ and $c\leq q$. By assumption, $c\vee p$ has an OA-factorization since $p$ is principal by Lemma 4.8. Since $\dim(L)\leq 2$, we have that $q$ is minimal over $c\vee p$. Let $c\vee p=\prod_{i=1}^{n}x_{i}$ where $x_{i}$ is an OA-element for each $i\in[1,n]$. We get that $x_{j}\leq q$ for some $j\in[1,n]$. Note that $m^{2}\nleq x_{j}$, and hence $x_{j}=q$. We infer that $c\vee p=q\ell$ for some $\ell\in L$. We conclude that $(c\vee p)m=c\vee p$ by the assumption, and thus $c\vee p=0$. Consequently, $c=p=0$, which contradicts the fact that $c\leq p$. Now assume that $q\neq qm$. Assume that $q$ is a nonminimal and nonmaximal prime element of $L$. Since $L$ is principally generated, there is a principal element $a\in L$ such that $a\leq q$ and $a\nleq qm$. Since $q$ is nonmaximal and $L$ is principally generated, there is a principal element $b\in L$ such that $b\leq m$ and $b\nleq q$. It remains to show that $x\leq a$ of each principal element $x\in L$ such that $x\leq q$. (Then $q=a$ is principal, since $L$ is principally generated and $a\leq q$.) Let $x\in L$ be principal such that $x\leq q$. Note that $xb^{2}$ is principal, and thus $a\vee xb^{2}$ is the join of two principal elements of $L$. Consequently, $a\vee xb^{2}$ has an OA-factorization. Since $a\vee xb^{2}\leq p$, there are $v,w\in L$ such that $v$ is an OA-element of $L$, $v\leq q$ and $a\vee xb^{2}=vw$. If $w\not=1$, then $w\leq m$, and hence $a\leq a\vee xb^{2}\leq qm$, a contradiction. Therefore, $a\vee xb^{2}=v$ is an OA-element. Assume that $b^{2}\leq a\vee xb^{2}$. Since $b^{2}$ is principal, we have that $1=(a\vee xb^{2}:b^{2})=(a:b^{2})\vee x$, and thus $(a:b^{2})=1$, since $L$ is quasi-local. Consequently, $b^{2}\leq a\leq q$, and hence $b\leq q$, a contradiction. This implies that $b^{2}\nleq a\vee xb^{2}$. Since $xb^{2}\leq a\vee xb^{2}$ and $a\vee xb^{2}$ is an OA-element, we conclude that $x\leq a\vee xb^{2}$. By [3, Theorem 1.4], it follows that $x\leq a$. ∎
Proposition 4.10.
Let $L$ be a principally generated C-lattice. If the join of any two principal elements has an OA-factorization, then $\dim(L)\leq 1$.
Proof.
First let every OA-element of $L$ be a prime element. We infer that $L$ is a ZPI-lattice by [15, Theorem 8]. Therefore, $\dim(L)\leq 1$ by [4, Theorem 2.6]. Now let there be an OA-element that is not a prime element. Then $L$ is quasi-local by Proposition 2.7. Let $m$ be the maximal element of $L$. We conclude by Proposition 2.8 that $m\neq m^{2}$. There exists some principal element $c\in L$ such that $c\nleq m^{2}$ and $c\leq m$.
Claim: $p\leq m^{2}$ for each nonmaximal prime element $p\in L$.
Let $p\in L$ be a nonmaximal prime element. By Lemma 4.9, we have that $\dim(L)\leq 2$. Therefore, we can assume without restriction that there are no prime elements of $L$ that are properly between $p$ and $m$ (i.e., for each prime element $r\in L$ with $p\leq r\leq m$, it follows that $r\in\{p,m\}$).
Next we show that $p=\bigwedge\{p\vee a\mid a\in L$ is principal and $a\nleq p\}$. Assume to the contrary that $p\not=\bigwedge\{p\vee a\mid a\in L$ is principal and $a\nleq p\}$. Then there is a principal element $y\in L$ such that $y\nleq p$ and $y\leq\bigwedge\{p\vee a\mid a\in L$ is principal and $a\nleq p\}$. Since $p<m$, there is a principal element $b\in L$ such that $b\nleq p$ and $b\leq m$. Since $by\in L$ is principal and $by\nleq p$, we have that $y\leq\bigwedge\{p\vee a\mid a\in L$ is principal and $a\nleq p\}\leq p\vee by$. It follows from [3, Theorem 1.4] that $y\leq p$, a contradiction. Therefore, $p=\bigwedge\{p\vee a\mid a\in L$ is principal and $a\nleq p\}$.
Assume that $p\nleq m^{2}$. It is sufficient to show that $m^{2}\leq p\vee a$ for each proper principal $a\in L$ with $a\nleq p$. (Then $m^{2}\leq\bigwedge\{p\vee d\mid d\in L$ is principal and $d\nleq p\}=p$, and hence $p=m$, a contradiction.) Let $a\in L$ be a proper principal element such that $a\nleq p$. Since $p$ is principal by Lemma 4.9, it follows that $p\vee a$ has an OA-factorization in $L$. Since $p\vee a\nleq m^{2}$ and $p\vee a$ is proper, we obtain that $p\vee a$ is an OA-element. Clearly, $p\vee a$ is not a nonmaximal prime element, and hence $m^{2}\leq p\vee a$ by Proposition 2.8. Consequently, $q\leq m^{2}$.∎(Claim)
It is sufficient to show that $r=0$ for each nonmaximal prime element $r\in L$. Let $r\in L$ be a nonmaximal prime element. Since $c$ has an OA-factorization and $c\nleq m^{2}$, we infer that $c$ is an OA-element of $L$. By the claim, it follows that $c$ is not a nonmaximal prime element of $L$. Observe that $r\leq m^{2}\leq c$ by the claim and by Proposition 2.8. Since $c$ is a (weak meet) principal element of $L$, $r$ is a prime element of $L$ and $c\nleq r$, we conclude that $r=cr$. Therefore, $r=0$ by [3, Theorem 1.4]. ∎
Proposition 4.11.
Let $(L,m)$ be a quasi-local principally generated C-lattice. If the join of any two principal elements of $L$ has an OA-factorization and $\dim(L)=1$, then $L$ is a domain.
Proof.
[…]
Proposition 4.12.
Let $L$ be a principally generated C-lattice and set $m={\rm J}(L)$. If the join of any two principal elements of $L$ has an OA-factorization, then $L$ satisfies one of the following conditions.
  1. (a)
    $L$ is a ZPI-lattice.
  2. (b)
    $L$ is a quasi-local lattice, $m^{2}$ is comparable and $m$ is a nilpotent element.
  3. (c)
    $L$ is a quasi-local lattice domain, $m^{2}$ is comparable and $\bigwedge_{n\in\mathbb{N}}m^{n}=0$.
Proof.
Let the join of any two principal elements of $L$ has an OA-factorization. Assume that every OA-element of $L$ is prime. By [15, Theorem 8], it follows that $L$ is a ZPI-lattice. Now, assume that there is an OA-element which is not a prime. We conclude that $(L,m)$ is a quasi-local C-lattice with $m^{2}\neq m$ by Propositions 2.7 and 2.8. Then there is a nonzero principal element $x\nleq m^{2}$. Clearly, $x$ is not the product of more than one OA-element, and thus $x$ is an OA-element.
First let $\dim(L)=0$. We show that $m^{2}$ is comparable. Let $z\in L$ be proper such that $z\nleq m^{2}$. There is a principal element $a\in L$ with $a\leq z$ and $a\nleq m^{2}$. Note that $a$ is an OA-element which is not a prime element. We have that $m^{2}\leq a$ by Proposition 2.8, and thus $m^{2}\leq z$. Consequently, $m^{2}$ is comparable. Clearly, 00 has an OA-factorization. This implies that $m^{k}\leq 0$ for some positive integer $k$, and hence $m^{k}=0$.
Now let $\dim(L)=1$. We obtain that $(L,m)$ is a quasi-local lattice domain by Proposition 4.11. To verify that $\bigwedge_{n\in\mathbb{N}}m^{n}=0$, assume the contrary that $\bigwedge_{n\in\mathbb{N}}m^{n}\neq 0$. There is a nonzero principal element $x\in L$ with $x\leq\bigwedge_{n\in\mathbb{N}}m^{n}$. Since $x$ is a product of $k$ OA-elements of $L$, we conclude that $m^{2k}\leq x\leq m^{4k}\leq m^{2k}$. In particular, we get that $x=x^{2}$. Since $x$ is principal, we have that $1=x\vee(0:x)$. We infer that $x=1$, a contradiction. Therefore, $\bigwedge_{n\in\mathbb{N}}m^{n}=0$. Finally, we show that $m^{2}$ is comparable. Let $z\in L$ be proper such that $z\nleq m^{2}$. There is a principal element $a\in L$ with $a\leq z$ and $a\nleq m^{2}$. We get that $a$ is a nonzero OA-element. Therefore, $m^{2}\leq a$ (since $\dim(L)=1$), and thus $m^{2}\leq z$. ∎
Theorem 4.13.
Let $L$ be a principally generated C-lattice and set $m={\rm J}(L)$. The following statements are equivalent.
  1. (1)
    $L$ is an OAFL.
  2. (2)
    $L$ is a COAFL.
  3. (3)
    The join of any two principal elements of $L$ has an OA-factorization.
  4. (4)
    $L$ satisfies one of the following conditions.
    1. (a)
      $L$ is a ZPI-lattice.
    2. (b)
      $L$ is a quasi-local lattice, $m^{2}$ is comparable and $m$ is a nilpotent element.
    3. (c)
      $L$ is a quasi-local lattice domain, $m^{2}$ is comparable and $\bigwedge_{n\in\mathbb{N}}m^{n}=0$.
Proof.
(1) $\Rightarrow$ (2) This is obvious.
(2) $\Rightarrow$ (3) Note that every principal element is compact, and hence the join of each two principal elements is compact. The statement is now immediately clear.
(3) $\Rightarrow$ (4) This follows from Proposition 4.12.
(4) $\Rightarrow$ (1) If $L$ is a ZPI-lattice, then clearly $L$ is an OAFL. Now let $L$ be not a ZPI-lattice. It is an immediate consequence of Lemma 4.4 that $L$ is an OAFL. ∎
Theorem 4.14.
Let $L$ be a principally generated C-lattice. The following statements are equivalent.
  1. (1)
    $L$ is a ZPI-lattice.
  2. (2)
    $L$ is a Prüfer OAFL.
  3. (3)
    $L$ is a Prüfer POAFL.
Proof.
(1$\Rightarrow$ (2) $\Rightarrow$ (3) This follows from [15, Theorem 8].
(3) $\Rightarrow$ (1) Let $L$ be a Prüfer POAFL. If $L$ is not quasi-local, then the prime elements coincide with the OA-elements. By Theorem 4.13, we infer that $L$ is a ZPI-lattice. Assume that $(L,m)$ is a quasi-local lattice with maximal element $m$. Then $m^{2}$ is comparable by Theorem 4.13. We know from Proposition 2.12(2) that each OA-element is either prime or equal to $m^{2}$. Therefore, $L$ is a ZPI-lattice. ∎
Finally, we provide a theorem that connects the various types of factorization lattices for a quasi-local principally generated C-lattice domain.
Theorem 4.15.
Let $(L,m)$ be a quasi-local principally generated C-lattice domain. The following statements are equivalent.
  1. (1)
    $L$ is an OAFL.
  2. (2)
    $L$ is a TAFL.
  3. (3)
    $L$ is a COAFL.
  4. (4)
    $L$ is a CTAFL that satisfies the ascending chain condition on prime elements.
  5. (5)
    $\dim(L)\leq 1$ and $L$ is a POAFL.
  6. (6)
    $\dim(L)\leq 1$ and $L$ is a PTAFL.
  7. (7)
    $\dim(L)\leq 1$, $m^{2}$ is comparable and $\bigwedge_{n\in\mathbb{N}}m^{n}=0$.
Proof.
(1) $\Leftrightarrow$ (3) $\Leftrightarrow$ (7) This follows from Theorem 4.13.
(1) $\Rightarrow$ (5) $\Rightarrow$ (6) This follows from Corollary 4.3.
(2) $\Leftrightarrow$ (6) $\Leftrightarrow$ (7) This is an immediate consequence of Theorem 3.11 and Proposition 3.12.
(2) $\Rightarrow$ (4) Clearly, $L$ is a CTAFL. Moreover, $\dim(L)\leq 1$ by Proposition 3.4. It is clear now that $L$ satisfies the ascending chain condition on prime elements.
(4) $\Rightarrow$ (6) Obviously, $L$ is a PTAFL. We infer by Proposition 3.7 that $\dim(L)\leq 1$. ∎
ACKNOWLEDGEMENTS. The first-named author was supported by the Austrian Science Fund FWF, Project Number P36742-N. The second-named author was supported by Scientific and Technological Research Council of Turkey (TUBITAK) 2219-International Postdoctoral Research Fellowship Program for Turkish Citizens, (Grant No: 1059B192202920).